OpenStudy (gabylovesyou):

The regular price of a child's entry ticket to a water park is $8 less than that for an adult's. The park offers half off of all entry tickets during the off-peak season. The Sandlers paid a total of $194 for 1 adult ticket and 4 child's tickets to the water park during the off-peak season. The following equation represents this situation, where x represents the regular price of an adult ticket. 194 = 1/2x + 2(x - 8) What is the regular price of a child's ticket? $68 $76 $84 $89

OpenStudy (gabylovesyou):

@dpaInc

OpenStudy (gabylovesyou):

Im going to close the question but i still need help lol

OpenStudy (anonymous):

let A = price of adult tickets so K = A-8 is the price of kid tickets so 194 = 1*A/2 + 4*K/2 \(\large 194=\frac{A}{2}+4\cdot \frac{K}{2} \) \(\large 194=\frac{A}{2}+2K \) now replacing K with A-8 we have: \(\large 194=\frac{A}{2}+2(A-8) \) \(\large 194=\frac{A}{2}+2A-16 \) \(\large 388=A+4A-32 \) multiplied both sides by 2 to clear the fraction \(\large 388=5A-32 \) \(\large 420=5A \) 84 = A so K = 84 - 8 = 76

OpenStudy (gabylovesyou):

so its B?? thanks :)